Find meassures of the each angle of the triangle
(2x+4)
(2x-9)
x

Answer by nerdybill(5404)

Remember:
The sum of the three interior angles of any triangle equals 180.
.
The problem defined the three angles as:
(2x+4) angle 1
(2x-9) angle 2
x angle 3
.
Our equation is then
(2x+4) + (2x-9) + x = 180
Grouping like-terms:
(2x+2x+x) + (4-9) = 180
5x + (-5) = 180
5x - 5 = 180
5x = 185
x = 37 degrees (angle 3)
.
For angle 2, we said it was:
(2x-9)
substituting our value of x:
(2(37)-9)
74-9
65 degrees (angle 2)
.
For angle 1, we said it was:
(2x+4)
substituting our value of x:
(2(37)+4)
74 + 4
78 degrees (angle 1)